Circular No. 6572

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 E. Costa and M. Feroci, Istituto Astrofisica Spaziale, CNR, Frascati; F.

Frontera and G. Zavattini, Universitá di Ferrara; L. Nicastro and E. Palazzi, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna; G. Spoliti, L. Di Ciolo, A. Coletta, G. D’Andreta, and J. M. Muller, BeppoSAX, Rome; and R. Jager, J. Heise, and J. in ’t Zand, Space Research Organization of the Netherlands, Utrecht, communicate: “The BeppoSAX Gamma Ray Burst Monitor (GRBM) was triggered on Feb. 28.123620 UT. A quick-look analysis of the GRBM light curve shows a moderately intense, multi-peaked γ-ray burst lasting ∼ 80 s, with one major structured peak followed by three minor peaks. The first peak intensity at maximum is ∼ 3700 counts/s, while the trailing peaks show a decreasing intensity from 600 to 150 counts/s. The γ-ray burst was detected by the Wide Field Camera (WFC) number 1 with a similar time profile and a peak countrate of ∼ 400 counts/s. Taking into account the off-axis position, the peak flux corresponds to ∼ 0.23 Crab. From the WFC image, we have derived a preliminary position of α = 5^{h}01^{m}57^{s}, δ = +11^{o}46′.4 (equinox 2000.0), with a 3′ error radius. A BeppoSAX TOO pointing of the Narrow Field Instruments was performed starting on Feb. 28.458, just 8 hr after the GRB trigger time. The relevant data analysis is in progress. Observations in all bands are solicited.”

SUPERNOVA 1997aa IN IC 2102 W.-d. Li, Y.-l. Qiu, Q.-y. Qiaom and J.-y. Hu, Beijing Astronomical

Observatory (BAO), on behalf of the BAO supernova survey, report the discovery of a supernova on CCD images of IC 2102 taken on Mar. 1.5 UT with the BAO 0.60-m reflector. SN 1997aa is located at α = 4^{h}51^{m}53^{s}.6, δ = −4^{o}57′36′′ (equinox 2000.0), which is 25′′ west and 29′′ south of the center of IC 2102. Available unfiltered CCD magnitudes: Mar. 1.5, ∼ 17.0; 2.5, ∼ 16.7. A low-resolution spectrum (40 nm/mm, range 350–930 nm) obtained by Q.-r. Yuan (Nanjing Normal University) and W. Liu and W.-d. Li (BAO) on Mar. 2.5 with the BAO 2.16-m telescope at Xinglong station shows that this is an type-II supernova prior to optical maximum. There are weak, yet conspicuous, H Balmer lines superimposed on a very blue continuum. The expansion velocity of the photosphere is 15 000 km/s (as derived from the Hα line). Hβ could be seen only marginally.

1997 March 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green